Please answer the question in full. Also, provide the work to show how got the answer.3. Beans Corp just paid a dividend of $4.00. The dividend will grow at a rate of 15% for 2 years and then settle down to a mo
Answer #1

D1=(4*1.15)=4.6

D2=(4.6*1.15)=5.29

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(5.29*1.06)/(0.09-0.06)

=186.913333

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.6/1.09+5.29/1.09^2+186.913333/1.09^2

=$165.99(Approx).